### v3

@@ 2.2 Computational Complexity
We will crucially use the following observation throughout the paper.
Observation 2.2.
The product of $O(\log n)$-many integers each of $O(\log n)$ bits can be computed in $\mathsf{AC^{0}}$.
Proof.
By [HAM02], the iterated multiplication of $k$ $k$-bit integers is in $\textsf{TC}^{0}$. In our setting, $k=m\in O(\log n)$. Thus, each Majority gate can be implemented using an AC circuit of constant depth and size $2^{O(\log n)}=\operatorname{poly}(n)$. The result now follows. ∎
@@
Time and space-restricted Turing machines. When considering complexity classes defined by Turing machines with a time bound $t(n)\in o(n)$, we use Turing machine with random access and a separate address (or index) tape. After writing an address, the machine can go to a query state reading the symbol from the input at the location specified by the address tape.
For functions $t(n),s(n)\in\Omega(\log n)$, the classes $\mathsf{DTISP}(t(n),s(n))$ and $\mathsf{NTISP}(t(n),s(n))$ are defined by deterministic (resp. nondeterministic) $t(n)$ time and $s(n)$ space bounded Turing machines. Note that there must be one Turing machine that simultaneously satisfies the time and space bound. For details we refer to [Vol99, Section 2.6]. For further reading on the connection to $\mathsf{quasiAC^{0}}$, we refer to [Bar92, FGSTT20].
Remark 2.3.
Because $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ is such a small class, when we say some function is computable in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, there is some ambiguity as to what this might mean (e.g., is there a separate output tape? Does it output one bit at a time?). Throughout this paper, where relevant (e.g., Lemma 3.1), it means that at the end of the computation, the result is written on the work tape of the Turing machine (and, hence, also is subject to the space bound).
@@
$\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{NTIME}(\mathrm{polylog}(n))\subseteq\mathsf{quasiAC^{0}}$.
Proof sketch.
Take the OR over all $2^{\mathrm{polylog}(n)}$ possible computation histories, of the $\mathsf{AC}^{0}$ circuit that verifies a computation history (the latter as in the proof of the Cook–Levin Theorem). ∎
@@
Let $L\in\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$, and suppose $L^{\prime}\leq_{dtt}^{\mathsf{AC^{0}}}$ via $g$. We show that $L^{\prime}$ is also in $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$. Let $V\in\mathsf{AC^{0}}(\mathcal{C})$ be a predicate such that $x\in L\iff[\exists^{\log^{2}|x|}w]\,V(x,w)$ for all strings $x$. Then we have $x\in L^{\prime}$ iff $\exists\,i\in\{1,\dots,k\},\exists\,w\in\{0,1\}^{\log^{2}n}:V(g(x)_{i},w)$, where $g(x)_{i}$ denotes the $i$-th string output by $g(x)=(y_{1},\dotsc,y_{k})$. Since $g$ is, in particular, polynomial size, we have $k\leq\operatorname{poly}(|x|)$, so we have the bit-length of the index $i$ is at most $O(\log|x|)$. Finally, since $\mathsf{AC^{0}}(\mathcal{C})$ denotes the oracle class, it is closed under $\mathsf{AC^{0}}$ reductions, and thus the function $(x,w,i)\mapsto V(g(x)_{i},w)$ is in $\mathsf{AC^{0}}(\mathcal{C})$, and therefore $L^{\prime}\in\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$. ∎
@@ 3 Order Finding and Applications
The following problem is in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$: On input of a multiplication table of a semi-group $S$, an element $s\in S$, and a unary or binary number $k\in\mathbb{N}$ with $k\leq\left|\mathinner{S}\right|$, compute $s^{k}$ (NB: Remark 2.3).
@@
On input of a multiplication table of a group $G$, an element $g\in G$, and $k\in\mathbb{N}$, we may decide whether $\operatorname{ord}(g)=k$ in $\forall^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. The same applies if, instead of $k$, another group element $h\in G$ is given with the question whether $\operatorname{ord}(g)=\operatorname{ord}(h)$.
@@
First check whether $g^{k}=1$ using Lem. 3.1. If yes (and $k>1$), we use $O(\log n)$ universally quantified co-nondeterministic bits to verify that for all $1\leq i<k$ that $g^{i}\neq 1$ using again Lem. 3.1. For the second form of input observe that $\operatorname{ord}(g)=\operatorname{ord}(h)$ if and only if for all $i\leq\left|\mathinner{G}\right|$ we have $g^{i}=1$ if and only if $h^{i}=1$. ∎
@@ 3.1 Application to isomorphism testing
Let $G$ be a finite simple group and $H$ be arbitrary. We can decide isomorphism between $G$ and $H$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
@@
As $G$ is a finite simple group, $G$ is determined up to isomorphism by (i) $|G|$, and (ii) the set of orders of elements of $G$: $\text{spec}(G):=\{\operatorname{ord}(g):g\in G\}$ [VGM09]. We may check whether $|G|=|H|$ in $\mathsf{AC}^{0}$. By Cor. 3.2, we may compute and compare $\text{spec}(G)$ and $\text{spec}(H)$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. The result now follows. ∎
@@
A finite group $G$ is nilpotent if and only if it is the direct product of its Sylow subgroups. Thus, for each prime $p$ dividing $|G|$, we identify the elements $X_{p}:=\{g:\operatorname{ord}(g)\text{ is a power of }p\}$ and then test whether $X_{p}$ forms a group. By Cor. 3.2, we can identify $X_{p}$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Verifying the group axioms is $\mathsf{AC^{0}}$-computable. The result follows. ∎
@@ 3.2 Application to membership testing
Proof sketch.
This follows by guessing suitable exponents and using Lem. 3.1 to compute the respective powers. ∎
@@
4 Quasigroup Isomorphism
In this section, we establish the following.
Theorem 4.1.
Quasigroup Isomorphism belongs to $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$.
Proof of Thm. 4.1.
Let $G$ and $H$ be quasigroups given as their multiplication tables. We assume that the elements of the quasigroups are indexed by integers $1,\dots,\left|\mathinner{G}\right|$. If $\left|\mathinner{G}\right|\neq\left|\mathinner{H}\right|$ (this can be tested in $\mathsf{DTIME}(\log n)$ by a standard binary search), we know that $G$ and $H$ are not isomorphic. Otherwise, let us write $n=\left|\mathinner{G}\right|$ and $k=\left\lceil\mathinner{2\log n}\right\rceil+1$.
  • that these sequences are actually cube generating sequences,
  • that $g_{i}\mapsto h_{i}$ induces an isomorphism.
Let us describe the first point for $G$ (for $H$ this follows exactly the same way): we universally verify for every element $g\in G$ (which can be encoded using $O(\log n)$ bits, hence, $\forall^{\log n}$) that we can existentially guess a sequence $(e_{1},\dots,e_{k})\in\{0,1\}^{k}$ (i.e. $\exists^{\log n}$) such that $g=P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$. We can compute this product in $\mathsf{DTISP}(\log^{3}n,\log n)$ by multiplying from left to right. [^8] Here, we could impose the additional requirement that the length of each row/column of the multiplication table is padded up to a power of two in order to get a bound of $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\log^{2}n,\log n)$) Each multiplication can be done in time $O(\log^{2}n)$ because we simply need to compute $i+j\cdot n$ for two addresses $i,j$ of quasigroup elements, write the result on the index tape and then read the corresponding product of group elements from the multiplication table. Moreover, note that for this procedure we only need to store one intermediate result on the working tape at any time. Thus, computing the product $P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$ can be done in time $O(\log^{3}n)$ and space $O(\log n)$ and it can be checked whether the result is $g$.
Again the products can be computed in time $O(\operatorname{poly}\log n)$ and space $O(\log n)$ as outlined above. Now, it remains to observe that we can combine the two $\forall^{\log n}$ blocks into one $\forall^{\log n}$ block because the check whether $g_{i}\mapsto h_{i}$ induces an isomorphism does not depend on the existentially guessed bits in the first check. ∎
We obtain several corollaries.
Corollary 4.2.
Isomorphism testing of two Steiner triple systems is in $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
Proof.
Given a Steiner triple system, we may write down a quasigroup $Q$ in the following manner. For each block $\{a,b,c\}$ in the Steiner triple system, we include the products $ab=c,ba=c,ac=b,ca=b,bc=a,cb=a$. Thus, we may write down the multiplication table for $Q$ in $\mathsf{AC^{0}}$. The Steiner triple system determines $Q$ up to isomorphism. The result now follows. ∎
Bose [Bos63] previously showed that pseudo-STS graphs with $>67$ vertices are STS graphs. Now given a block-incidence graph from a Steiner $2$-design with bounded block size, we can recover the underlying design in $\mathsf{AC^{0}}$ [Lev23, Proposition 4.7]. This yields the following.
Corollary 4.3.
Isomorphism testing of pseudo-STS graphs belongs to $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
As isomorphism testing of Steiner $(t,t+1)$-designs is $\exists^{\log^{2}n}\mathsf{AC^{0}}$-reducible to isomorphism testing of Steiner triple systems [BW13] (cf., [Lev23, Corollary 4.11]), we obtain the following corollary.
Corollary 4.4.
Isomorphism testing of Steiner $(t,t+1)$-designs is in $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
Given two Latin square graphs $G_{1},G_{2}$, we may recover corresponding Latin squares $L_{1},L_{2}$ in $\mathsf{AC^{0}}$ (cf. [Lev23, Lemma 3.9]). Now $G_{1}\cong G_{2}$ if and only if $L_{1}$ and $L_{2}$ are main class isotopic [Mil78, Lemma 3]. We may decide whether $L_{1},L_{2}$ are main class isotopic using an $\mathsf{AC^{0}}$-computable disjunctive truth-table reduction to Quasigroup Isomorphism (cf. [Lev23, Remark 1.6]). As $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ is closed under $\mathsf{AC^{0}}$-computable dtt reductions (Fact 2.7), we get the following corollary.
Corollary 4.5.
Isomorphism testing of Latin square graphs belongs to $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
Remark 4.6.
We note that the reduction outlined in [Mil78, Theorem 2] and [Lev23, Remark 1.6] in fact allow us to determine whether two quasigroup are isotopic, and not just main class isotopic. Thus, we have an $\mathsf{AC^{0}}$-computable disjunctive truth-table reduction from Latin Square Isotopy to Quasigroup Isomorphism. We now obtain the following corollary.
Corollary 4.7.
Latin Square Isotopy belongs to $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.

### v4

@@ 2.2 Computational Complexity
We will crucially use the following throughout the paper.
Theorem 2.2 ([HAM02, Theorem 5.1]).
The product of $(\log n)^{O(1)}$-many integers each of $(\log n)^{O(1)}$ bits can be computed in $\mathsf{AC^{0}}$.
@@
Time and space-restricted Turing machines. When considering complexity classes defined by Turing machines with a time bound $t(n)\in o(n)$, we use Turing machines with random access and a separate address (or index) tape. After writing an address, the machine can go to a query state reading the symbol from the input at the location specified by the address tape. As usual, the machines are otherwise allowed to have multiple work tapes.
For functions $t(n),s(n)\in\Omega(\log n)$, the classes $\mathsf{DTISP}(t(n),s(n))$ and $\mathsf{NTISP}(t(n),s(n))$ are defined to consist of decision problems computable by deterministic (resp. nondeterministic) $t(n)$ time and $s(n)$ space bounded Turing machines. Be aware that there must be one Turing machine that simultaneously satisfies the time and space bound. For details we refer to [Vol99, Section 2.6]. For further reading on the connection to $\mathsf{quasiAC^{0}}$, we refer to [Bar92, FGSTT20].
We frequently use $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ and $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. However, because of how small the time and space bounds are for these classes, when we abuse notation to say some function (not necessarily decision problem) is computable in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, there is some ambiguity as to what this might mean. We use the following two equivalent definitions.
[…]
Definition/Lemma 2.3.
For any function $f\colon\{0,1\}^{*}\to\{0,1\}^{*}$ with $|f(x)|\leq O(\log|x|)$, the following are equivalent:
  1. 1.
    $f$ is computable by a Turing machine, using $\operatorname{poly}(\log n)$ time and $O(\log n)$ space, that halts with the result written in a specified place on its work tape.
  2. 2.
[…]
In either case, we say $f$ is computable in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$.
Proof.
($\Rightarrow$) Suppose $f$ is computable by a Turing machine $M$ using $\operatorname{poly}(\log n)$ time and $O(\log n)$ workspace (including the output). Then the following machine $N$ computes the function $(x,i)\mapsto f(x)_{i}$. On input $x,i$, $N$ simulates $M(x)$ while leaving $i$ untouched. After the simulation has completed, one of $N$’s worktapes contains $f(x)_{i}$, while $i$ is still on the input tape. Next, using another worktape to keep a counter, $N$ walks through $f(x)$ until it gets to the $i$-th bit, and outputs that bit of $f(x)$.
The entirety of the computation after the simulation of $N$ uses an amount of space that is $|i|\leq O(|f(x)|)\leq O(\log|x|)$, and time that is at most $|i|^{2}\leq O(\log^{2}|x|)$ (in fact, by noting that incrementing a counter by $+1$ $k$ many times only incurs $O(k)$ many changes to the bits of the counter, this can be reduced to $O(\log|x|)$ as well).
($\Leftarrow$) Suppose the function $(x,i)\mapsto f(x)_{i}$ is computable by a Turing machine $N$ using $\operatorname{poly}(\log n)=\operatorname{poly}(\log|x|)$ time and $O(\log n)=O(\log|x|)$ space. The following machine $M$ computes $f$ in the manner in which, when started with $x$ on its input tape, at the end of the computation $f(x)$ is the only thing written on the work tape. On input $x$, $M$ simulates $N(x,0),N(x,1),\dotsc$, writing those bits in the specified place on its work tape, until it reaches an $i$ such that $N(x,i)=\bot$. As $N$ was called $|f(x)|+1$ times, this multiplies the time complexity of $N$ by $|f(x)|\leq O(\log|x|)$. However, the space need does not increase except to keep track of the integer $i$ for which the machine will next simulate $N(x,i)$, plus the space for the output, as the space to simulate $N(x,i)$ can be re-used when simulating $N(x,i+1)$. Thus the space used is that of $N$, plus the space for the output, plus the space to maintain $i$, which is $|i|\leq\log(|f(x)|+1)\leq O(\log\log|x|)$, so the total space is still $O(\log|x|+\log\log|x|)=O(\log|x|)$. The additional clearing and marking at the end at most doubles the time and adds a constant amount of space usage. ∎
@@
$\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{NTIME}(\mathrm{polylog}(n))\subseteq\mathsf{quasiAC^{0}}$. In particular, any decision problem in $\mathsf{NTIME}(\mathrm{polylog}(n))$ is computable by a quasi-polynomial-size DNF (a particular kind of depth-2 $\mathsf{quasiAC^{0}}$ circuit).
Furthermore, decision problems in $\mathsf{DTIME}(\mathrm{polylog}(n))$ are computable by quasi-polynomial-size CNFs as well.
Proof.
Given a $\mathrm{polylog}(n)$-time nondeterministic Turing machine $M$, we get a decision tree $T_{M}$ that decides that same language as follows. Each node of the tree corresponds to any time $M$ either queries an input bit or makes a nondeterministic guess. The two children of such a node correspond to whether the queried (resp., nondeterministically guessed) bit was 0 or 1. The leaves of the tree are labeled accepting or rejecting according to whether $M$ accepts or rejects at that point (note: once no more input bits are queried and no more nondeterministic guesses are made, the answer is fully determined, even though $M$ may deterministically spend an additional $\mathrm{polylog}(n)$ steps figuring out what the answer should be). The decision tree has height at most $\mathrm{polylog}(n)$, and thus at most quasi-polynomially many branches. We get a DNF by taking the disjunction over all accepting paths of the conjunction of the answers to the input queries on those paths.
If there are no non-deterministic guesses made, i.e. our machine $M$ was deterministic, then we can get a CNF from the above decision tree by using De Morgan’s law: negate the leaves of the decision tree, get the corresponding DNF of the negated tree, then negate the DNF. (This does not work in the nondeterminstic case, as the disjunction includes all inputs such that there exists nondeterministic guesses that lead to an accepting leaf, and when we negate the leaves we end up with co-nondeterminism instead of non-determinism.) ∎
@@
Let $L\in\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$, and suppose $L^{\prime}\leq_{dtt}^{\mathsf{AC^{0}}}L$ via $g$. We show that $L^{\prime}$ is also in $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$. Let $V\in\mathsf{AC^{0}}(\mathcal{C})$ be a predicate such that $x\in L\iff[\exists^{\log^{2}|x|}w]\,V(x,w)$ for all strings $x$. Then we have $x\in L^{\prime}$ iff $\exists\,i\in\{1,\dots,k\},\exists\,w\in\{0,1\}^{\log^{2}n}:V(g(x)_{i},w)$, where $g(x)_{i}$ denotes the $i$-th string output by $g(x)=(y_{1},\dotsc,y_{k})$. Since $g$ is, in particular, polynomial size, we have $k\leq\operatorname{poly}(|x|)$, so we have the bit-length of the index $i$ is at most $O(\log|x|)$. Finally, since $\mathsf{AC^{0}}(\mathcal{C})$ denotes the oracle class, it is closed under $\mathsf{AC^{0}}$ reductions, and thus the function $(x,w,i)\mapsto V(g(x)_{i},w)$ is in $\mathsf{AC^{0}}(\mathcal{C})$, and therefore $L^{\prime}\in\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathcal{C})$. ∎
@@ 3 Order Finding and Applications
The following problem is in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$: On input of a multiplication table of a semigroup $S$, an element $s\in S$, and a unary or binary number $k\in\mathbb{N}$ with $k\leq\left|\mathinner{S}\right|$, compute $s^{k}$ (NB: Definition/Lemma 2.3).
Note that, because the multiplication table is part of the input, the input size $n$ is always at least $|S|$.
@@
On input of a multiplication table of a group $G$, an element $g\in G$, and $k\in\mathbb{N}$, we may decide whether $\operatorname{ord}(g)=k$ in $\forall^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. The same applies if, instead of $k$, another group element $h\in G$ is given with the question whether $\operatorname{ord}(g)=\operatorname{ord}(h)$. In particular, both questions can be decided by quasi-polynomial-size CNFs (a particular case of depth-2 $\mathsf{quasiAC^{0}}$).
@@
First check whether $g^{k}=1$ using Lem. 3.1. If yes (and $k>1$), we use $O(\log n)$ universally quantified co-nondeterministic bits to verify that for all $1\leq i<k$ that $g^{i}\neq 1$ using again Lem. 3.1. For the second form of input observe that $\operatorname{ord}(g)=\operatorname{ord}(h)$ if and only if for all $i\leq\left|\mathinner{G}\right|$ we have $g^{i}=1$ if and only if $h^{i}=1$.
For the depth-2 upper bound: we can compute the decision problems $(g,i)\mapsto[g^{i}=1]$, $(g,i)\mapsto[g^{i}\neq 1]$, and $(g,h,i)\mapsto[g^{i}=1\leftrightarrow h^{i}=1]$ in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which in turn can be decided by quasi-polynomial-size CNFs. Implementing the above strategy, the $\forall^{\log n}$ becomes an AND gate of fan-in $\operatorname{poly}(n)$. That can be merged with the AND gates at the top of the aforementioned CNFs to get a single CNF of quasi-polynomial size. ∎
@@ 3.1 Application to isomorphism testing
Let $G$ be a finite simple group and $H$ be arbitrary. We can decide isomorphism between $G$ and $H$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$
@@
As $G$ is a finite simple group, $G$ is determined up to isomorphism by (i) $|G|$, and (ii) the set of orders of elements of $G$: $\text{spec}(G):=\{\operatorname{ord}(g):g\in G\}$ [VGM09]. We may check whether $|G|=|H|$ in $\mathsf{AC}^{0}$. By Cor. 3.2, we may compute and compare $\text{spec}(G)$ and $\text{spec}(H)$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. ∎
@@
A finite group $G$ is nilpotent if and only if it is the direct product of its Sylow subgroups. As $G$ is given by its multiplication table, we may in $\mathsf{AC^{0}}$ compute the prime factors of $|G|$. Thus, for each prime $p$ dividing $|G|$, we identify the elements $X_{p}:=\{g:\operatorname{ord}(g)\text{ is a power of }p\}$ and then test whether $X_{p}$ forms a group. By Cor. 3.2, we can identify $X_{p}$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. Verifying the group axioms is $\mathsf{AC^{0}}$-computable. The result follows. ∎
@@ 3.2 Application to membership testing
Proof.
This follows by guessing suitable exponents and using Lem. 3.1 to compute the respective powers. In order to keep the space logarithmically bounded, we do not simply guess $g_{1},\dotsc,g_{m}\in X$—which could take up to $\log^{2}n$ bits—but rather guess the $g_{i}$ sequentially and only store the running product. That is, first guess $g_{1}$ and $e_{1}$ using $2\log n$ bits, and compute $g_{1}^{e_{1}}$. Inductively suppose we have computed $g_{1}^{e_{1}}\dotsb g_{i}^{e_{i}}$. Note that we do not need to store the elements $g_{1},\dotsc,g_{i}$ and exponents $e_{1},\dotsc,e_{i}$, we only keep the product $g_{1}^{e_{1}}\dotsb g_{i}^{e_{i}}$, which takes $\log n$ bits. Then we guess $g_{i+1}$ and $e_{i+1}$, and continue. ∎
@@
Proof.
Commutative semigroups have the $\log n$ power basis property [Fle22, Lem. 4.2]. ∎